1. A study shows that the average daily coffee consumption of a 20-30 years old students is 3 cups per day. A university claims that their students tend to drink less than 3 cups. They selected 20 students and found the mean of 3.5 with a standard deviation of 1.5 cups. Use 0.01 level of significance to test their claim.



Answer the following:



Parameter:


Claim:


Claim ( in symbol ):


Ho: Ho:


Ha: Ha:


What is the significance level or a?


Is it two-tailed or one-tailed test?

1
Expert's answer
2022-04-27T09:34:08-0400

Let "X=" the average daily coffee consumption.

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge3"

"H_1:\mu<3"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.01," "df=n-1=20-1=19" degrees of freedom, and the critical value for a right-tailed test is "t_c = -2.539483."

The rejection region for this left-tailed test is "R = \{t: t<-2.539483\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{3.5-3}{1.5/\sqrt{20}}\approx1.490712"

Since it is observed that "t =1.490712 >-2.539483=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for left-tailed, "df=19" degrees of freedom, "t=1.490712" is "p=0.923774," and since "p=0.923774>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is less than 3, at the "\alpha = 0.01" significance level.